OpenStudy (anonymous):

use elimination method to solve: a + 9b = 19 a + 2b = -2

OpenStudy (anonymous):

a + 9b = 19 (-1)(a + 2b = -2) a + 9b = 19 -a - 2b = 2 Cross out the positive "a" and the negative "a" 7b = 21 b = 3 Then plug 3 in for "b" for one of the equations. a + 9(3) = 19 a + 27 = 19 -27 -27 a = -8

OpenStudy (anonymous):

thank u!

OpenStudy (anonymous):

No problem